OpenStudy (anonymous):

Find the absolute value of complex number... |-2 + 10i|

OpenStudy (anonymous):

I remember a + b^2, so -2 + (-1) ^Because the i^2 makes the whole thing -1 right?

OpenStudy (anonymous):

Yet the final answer is \[2\sqrt{26}\]

OpenStudy (zarkon):

\[|a+bi|=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

Oh, square root of a^2 + b^2. Thanks.